What is the quadratic form for f(m)=(2m-5)²?
Answer by sdmmadam@yahoo.com(530)

f(m)=(2m-5)²
= [(2m)^2 + 2X(2m)X(-5) + (-5)^2]
(using (a+b)^2 = (a^2) +2(a)(b) + (b^2) and here a = 2m and b = (-5)
=(4m^2-20m +25)
which is the desired quadratic form